For each population, find the mean of the sampling distribution of the sample mean with an indicating sample size,n.


A.7,10,19,20 ; n=2


B.46,51,55,59,61 ; n=3


C.24,26,29,32,35 ; n=4

1
Expert's answer
2022-04-05T09:27:07-0400

A) sample distribution mean

7,10 8.5

7,19 13

7,20 13.5

10,19 14.5

10,20 15

19,20 19.5

mean μ= (8.5+ 13 + 13.5 +14.5 + 15+ 19.5)/6

=14

B) sample distribution mean

46,51,55 50.7

46,51,59 52

46,51,61 52.7

46,55,59 53.3

46,55,61 54

46,59,61 55.3

51,55,59 55

51,55,61 55.7

51, 59,61 57

55,59,61 58.3

the mean μ =(50.7 +52+52.7+53.3+54+55.3+55+55.7+57+58.3)/10

=54.4



C) sample distribution mean

24,26,29,32 27.75

24,26,29,35 28.5

24,26,32,35 29.25

24,29,32,35 30

the mean μ = (27.75+28.5+29.25+30)/4

=28.88